Somebody help me pls!. Been stuck with this problem.
A woman earns $800 in interest from two accounts in a year. She has four times as much invested at 7% as she does at 4%. Use elimination method to find out how much is invested in each account. Willl be grateful for any help you can provide.

Answer by scott8148(6628)

y = 4x __ y - 4x = 0

(7%)(y) + (4%)(x) = 800 ___ 7y + 4x = 80000

adding the two secondary equations will eliminate x